OpenStudy (anonymous):

How do you write 42% , 72% , 88% , and 35% as a fraction in simplest form.

OpenStudy (whpalmer4):

x% just means x/100. \( 33\% = \dfrac{33}{100}\) Once you have the fraction, make sure you can't divide out any common factors. For example, \(50\% = \dfrac{50}{100} = \dfrac{1}{2}\)

OpenStudy (campbell_st):

any percentage can be written as a fraction by writing it over 100 so 42% = 42/100 now you need to find common factors the divide evenly into both the numberator and denominator start with 2 21/50 no more common factors so 42% = 21/50 use this for the others...

OpenStudy (anonymous):

Im not really good at finding factors.i think i got 43% \[43/100 \] divide by one equals 43 equal 100

OpenStudy (anonymous):

@Princessaja is correct

OpenStudy (anonymous):

@soty2013 thanks